OpenStudy (anonymous):

Your friend tells you that at 1:00 pm the temperature outside was 67 degrees Fahrenheit and at 5:00 pm, that same day, the temperature was 65 degrees Fahrenheit. He also states that the average rate of change of the temperature between 1:00 pm and 5:00 pm increased by 2 degrees Fahrenheit. Determine if your friend is correct and explain why or why not.

OpenStudy (gudden):

Well the temperature... has first of all decrease.. not increased.. and your friend, is wrong in saying that very thing...

geerky42 (geerky42):

Ueah, temperature went from 67 to 65, hence rate of change cannot be increased

OpenStudy (anonymous):

thx

OpenStudy (harsha19111999):

Its wrong. Though your friend have said it has decreased by 2 degrees, it wrong then also

OpenStudy (harsha19111999):

Because you have asked average change= 65-67/5-1 = -2/4 = -1/2 = -0.5 is the average change rate per hour

OpenStudy (harsha19111999):

So the answer would be -0.5 F/hr

OpenStudy (anonymous):

its not per hour its from 1 to 5 i dont see them mentioning rate change per hour

geerky42 (geerky42):

Well, this doesn't really matter, friend is wrong because he used "increased" instead of "decreased."

OpenStudy (anonymous):

rate change is also per minute then per second,gives you multiple answers

OpenStudy (harsha19111999):

But we have here hours right

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

we have time here where you considered 4 hours ^^ which could be 240 mins,1440 secs :P

OpenStudy (harsha19111999):

For example, if you consider it as a function of time, you could have the time in minutes or seconds.But, the most likely used term is hours or the given time

OpenStudy (anonymous):

nooooo lol anyway

OpenStudy (harsha19111999):

lol. u give up?

OpenStudy (anonymous):

The question isn't specific enough to state how often the temperature was taken. It's clear it was taken more than once, average rate of change, but it could have been taken every 30min, every hour, every 5 min, etc. The answer is clear without needing to use numbers. It's a trick question where you think you need to solve an equation but don't.

OpenStudy (triciaal):

it's still -2 degrees over 4 hours for the average rate of change 1pm to 5pm = 4hours later you can convert to whatever unit you want

OpenStudy (anonymous):

yes I know, but this is my point. There's no need to find an answer due to lack of information. You (not you; a previous response) can't say the answer is -x°F/hr since the question isn't clear on what the unit of time would be used. it could be -x°/min for all we know.

OpenStudy (anonymous):

thats what am telling him :P he said even for decrease its wrong thats why i stated that argument

OpenStudy (triciaal):

remember the word average

OpenStudy (triciaal):

it still change by an average of -2 degrees over the 4 hour period

OpenStudy (anonymous):

yes I agree with that, -2° over those 4 hours. so it would be -.5° per hour. But the question is not asking what would the average rate of change be per hour. It's just asking if he's right or wrong.

OpenStudy (triciaal):

@Debs_ my point was to agree with you about the rate. Review the responses from the top. to answer the stated question the friend is wrong because the temperature showed a decrease over the only 2 readings given.

OpenStudy (anonymous):

@triciaal oh ok. I'm dumb. Please excuse me :P